# Adaptive Multimodal Person Recognition: A Robust Framework for Handling Missing Modalities

### v1

@@ Abstract
Person recognition systems often rely on audio, visual, or behavioral cues, but real-world conditions frequently result in missing or degraded modalities. To address this challenge, we propose a Trimodal person identification framework that integrates voice, face, and gesture modalities, while remaining robust to modality loss. Our approach leverages multi-task learning to process each modality independently, followed by a cross-attention and gated fusion mechanisms to facilitate interaction across modalities. Moreover, a confidence-weighted fusion strategy dynamically adapts to missing and low-quality data, ensuring optimal classification even in Unimodal or Bimodal scenarios. We evaluate our method on CANDOR, a newly introduced interview-based multimodal dataset, which we benchmark for the first time. Our results demonstrate that the proposed Trimodal system achieves 99.18% Top-1 accuracy on person identification tasks, outperforming conventional Unimodal and late-fusion approaches. In addition, we evaluate our model on the VoxCeleb1 dataset as a benchmark and reach 99.92% accuracy in Bimodal mode. Moreover, we show that our system maintains high accuracy even when one or two modalities are unavailable, making it a robust solution for real-world person recognition applications. The code and data for this work are publicly available  [^1] https://github.com/areffarhadi/Multimodal-Rec.
1 Introduction
@@
Typical person identification systems rely heavily on a single modality and risk severe performance degradation whenever that modality fails. This challenge calls for an approach that can effectively leverage the strengths of multiple modalities while downplaying those that are missing or degraded [fa2024comparative, fa2023facial]. However, Multimodal person identification, which combines cues from several modalities such as face, gait, gesture, and voice, can deliver robust and accurate recognition. Utilizing typical fusion methods [fa2024comparative], which concatenate all features, pose still some challenges in real-world conditions, such as one or more modalities frequently becoming unreliable due to occlusions, poor lighting, background noise, or short utterances [farhad2024analysis, Li_2024_CVPR].
@@
This work addresses concerns about the low performance of single-modality systems in out-of-domain conditions and the performance decrements of multimodal systems when they encounter modality loss. The proposed system introduces an adaptive fusion strategy that integrates face, gesture, and voice modalities within a unified neural architecture (more in Section 3). The core concept of our approach to handle modality loss are: confidence weighting, cross-modality bridging and multi-stage fusion with mistake correction. In confidence weighting each modality is equipped with a confidence network that quantifies the reliability of its current modality.
@@
The second key concept to handle modality is cross-modality bridging [farhadipour2025towards], which is realized through cross-attention [vaswani2017attention]. Using this strategy, we can combine a master modality with several auxiliary modalities in a learnable manner to create a more effective representation vector. Finally, we incorporate a multi-stage fusion approach that merges confidence-weighted logits with a gated fusion network, followed by a mistake-correction mechanism. This correction module rectifies overconfident errors by re-examining logit outputs from the face, gesture, and voice pathways alongside the fused ensemble. The result is a flexible pipeline capable of robust person identification even in substantial modality loss. To the best of our knowledge, our system addresses a relatively unexplored area of combining gestures with face and voice modalities.
@@
To validate our approach, we conducted experiments on the CANDOR dataset [reece2023CANDOR], a newly introduced interview-style dataset. Establishing CANDOR as a benchmark for person identification is also one of the primary objectives of this work. Furthermore, we used the audiovisual VoxCeleb1 dataset [nagrani2017voxceleb] as a benchmark to evaluate our approach compared to previous work. To this end, we utilized the original training, validation, and test splits provided by the dataset. It is essential to note that the multimodal version of VoxCeleb1 consists only of still images, which prevents capturing gesture modalities. Therefore, our system has been developed specifically for Bimodal using the VoxCeleb1 dataset.
@@
In the following sections, we delve deeper into the specifics of our model and experimental findings. First, we discuss related prior work on modality fusion and highlight how existing methods work with different modalities. Next, we present the proposed Trimodal system architecture. We then detail the partitioning of the CANDOR dataset and our data augmentation, which helps simulate real-world challenges. Our experimental results illustrate how this design yields robust identification under Unimodal, Bimodal, and fully Trimodal conditions.
2 Previous Work
@@
Initial face-gait fusion methods generally fused outputs at the score or decision level [Geng2008]. Some other early works relied on hand-crafted features and static fusion weights. For instance, Geng et al. proposed an adaptive scheme to combine face and gait cues in video, dynamically adjusting each modality’s contribution [Geng2008].
@@
Feature-level fusion then emerged, with approaches like subspace projection or concatenation to form a unified representation of face and gait [Xing2015]. While these multi-level fusion frameworks improved accuracy, they still relied on engineered features, limiting their performance in complex real-world scenarios.
@@
Studies commonly employ parallel pipelines for each modality, fusing their outputs at the feature or score level [Rahman2019, Maity2021]. Aung et al. employed transfer learning with Convolutional Neural Network-based (CNN) for both face and gait, concatenating feature vectors into a joint embedding and attaining above 97% accuracy [Aung2022]. Similarly, Manssor et al. tackled night-time surveillance by integrating face and gait recognition within a YOLO-based pipeline, highlighting the advantage of multimodal cues in adverse conditions [Manssor2021].
@@
Recent research focuses on attention-based and context-aware fusion, leveraging transformer-like mechanisms to dynamically weight each modality. Prakash et al. introduced Adapt-FuseNet with a keyless attention module that assigns importance to face or gait features based on reliability, achieving improved identification performance in video-based tasks [Prakash2023]. Zou and Wu proposed a robust hybrid framework with a dynamic weighting and distillation module that aligns and fuses face-gait features, improving recognition even under modality degradation [Zou2023]. These methods exemplify a shift from static to data-driven adaptive fusion, critical in unconstrained environments.
@@
While most deep fusion models rely on extensive labeled training data, emerging self-supervised techniques aim to reduce annotation dependency [Catruna2021]. Such methods exploit natural co-occurrence of face and gait in videos, using tasks like contrastive learning or cross-modal reconstruction to learn shared identity embeddings without explicit identity labels.
@@
Tiong et al. [Tiong_2024_CVPR] introduced a Transformer-based approach called Flexible Biometrics Recognition to fuse face and periocular features via dedicated attention modules and improved cross-modal interaction. Praveen and Alam [Praveen_2024_FG] addressed face-voice verification using recursive joint cross-attention. By applying cross-attention repeatedly, the model aligns voice and face embeddings at multiple stages, progressively refining a shared audio-visual representation.
@@
Fig. 1: Proposed Trimodal system. Consists of feeding each modality into its corresponding encoder and processing pathway. Subsequently, the outputs from these pathways are utilized within a Trimodal cross-attention block. Finally, all information is integrated into the decision-making block.
@@
Li et al. [Li_2024_CVPR] proposed a learnable multimodal tokenizer for RGB, infrared, sketches, and textual information that encodes each modality into a shared embedding space, which is then fed to a frozen transformer backbone. Crucially, the proposed model synthesized missing modalities during training, fostering robustness under incomplete inputs.
@@
Several works have reported speaker identification results on the VoxCeleb1 dataset, leveraging both supervised and self-supervised methods. Nagrani et al. proposed a cross-modal voice-to-face matching network and achieved 81.0% Top-1 accuracy on a shared speaker pool [nagrani2018seeing]. Yadav and Rai introduced a VGG-style CNN trained with Softmax and center loss, reaching 89.5% Top-1 accuracy [yadav2018learning]. Chung et al. proposed environment-adversarial training with a thin ResNet-34 regularized via confusion loss, achieving 89.0% Top-1 accuracy [chung2019delving].
@@
Shah et al. designed a dual-stream architecture fusing VGGFace and VGGVox embeddings, which reached 97.2% Top-1 accuracy in the bimodal setting [shah2023speaker]. Niizumi et al. introduced a self-supervised masked modeling duo approach that attained 81.2% Top-1 accuracy without any labeled fine-tuning [niizumi2024masked], while Cheng et al. proposed a pipeline for integrating Speech Anonymization and Identity Classification (SAIC) pipeline for joint speaker anonymization and classification, improving performance to 96.1% [cheng2024saic]. Anidjar et al. employed Wav2Vec2 embeddings combined with a lightweight CNN and aggressive data augmentation, resulting in 90.5% Top-1 accuracy [anidjar2024harnessing].
3 Trimodal Person Identification
Figure 1 provides an illustration of the proposed system. Each input video is represented by three feature vectors, extracted from three efficient pre-trained encoders and the beginning of modality-specific pathways. For the face modality, we utilized YOLOv8 [redmon2016you] to extract the face from the frames of a video, and MagFace [meng2021magface] for face quality assessment to determine the best facial frame per video. IR101-Adaface [deng2019arcface] was used for a 512-dimensional face embedding extraction. This model was pre-trained on 12 million facial images from the WebFace12M dataset. However, Timesformer [bertasius2021space] was used as the encoder for extracting 768-dimensional gesture embeddings based on 16 equally spaced frames from each video. This model was originally trained for action recognition, and we utilized the pretrained version. Given that the video lengths range between 4 and 16 seconds, the interval between the selected frames varies from 250 milliseconds to 1 second. Finally, a Resnet293 [lin2024voxblink2] pretrained on Voxblink2 and VoxCeleb1 datasets to extract a 256-dimensional voice embedding vector from each utterance.
3.1 Modality-Specific Pathways
@@
Each modality undergoes two main transformations, consisting of a fully connected block and a self-attention Block. Let $SA_{\text{face}}$ be the face pathway self-attention. Then, for the face modality, we can write:
@@
where $\mathbf{x}^{(\text{face})}\!\in\mathbb{R}^{512}$ is the output of two dense layers, and $c^{(\text{face})}\!\in(0,1)$ is a scalar confidence value obtained via a small feed-forward head and a Sigmoid $\sigma(\cdot)$ function. This design allows the network to both refine face features and dynamically estimate their reliability, facilitating confidence-weighted fusion in subsequent modules. Moreover, $\mathbf{p}^{(\text{face})}\in\mathbb{R}^{K}$ (with $K$ the number of classes) is the vector of unnormalised scores before softmax or logits produced by the face-specific classifier. Analogous expressions hold for other modalities.
3.2 Trimodal Cross-Attention
After $\mathbf{z}^{(\text{face})}$, $\mathbf{z}^{(\text{gest})}$, and $\mathbf{z}^{(\text{voice})}$ are extracted, they are fed into a cross-attention module in which all three modalities are jointly refined. This module bridges information across the face, gesture, and voice streams, ensuring that each modality interacts with the others rather than being processed in isolation.
This is achieved in two key steps, comprising Cross-Modality Projection and Cross-attention Processing. In cross-modality projection, each modality receives a projection of the other two modalities into its respective feature space.
The cross-attention module plays a key role in bridging the three modalities, ensuring that information is shared, refined, and learned adaptively. This component significantly improves the robustness and accuracy of person identification. For instance, the face cross sub-block aggregates gesture and voice embeddings into the face branch:
@@ 3.2 Trimodal Cross-Attention
3.3 Fusion Classification
Two main mechanisms are employed to combine the three modalities for final classification, consisting of gated feature fusion and confidence-weighted fusion.
The feature fusion module is designed to combine multimodal information while suppressing noise intelligently. It consists of feature concatenation that brings together face, gesture, and voice features into a unified vector.
$\mathbf{z}_{x}^{\text{concat}}=\left[\mathbf{z}_{x}^{(\text{face})},\mathbf{z}_{x}^{(\text{gest})},\mathbf{z}_{x}^{(\text{voice})}\right]\in\mathbb{R}^{960}$ (5)
The fusion gate acts as a dynamic weighting mechanism, selectively emphasizing useful information and filtering out irrelevant noise. A learnable gating mechanism determines how much each feature dimension contributes:
@@ 3.3 Fusion Classification
where $W_{1},W_{2}$ are learnable weight matrices, $\phi(\cdot)$ is a non-linear activation function, $\sigma(\cdot)$ is the Sigmoid function, ensuring gating values in $(0,1)$.
@@
$\mathbf{z}^{\text{fused}}=\mathbf{z}_{x}^{\text{concat}}\odot\mathbf{g},$ (7)
where $\odot$ represents element-wise multiplication. This means important features (high gate values) pass through strongly and irrelevant or noisy features (low gate values) are suppressed.
@@
The Confidence-Weighted Fusion mechanism ensures that the contribution of each modality is dynamically adjusted based on confidence scores. Each modality has an associated confidence score ($c^{(\text{face})},c^{(\text{gest})},c^{(\text{voice})}$) which are learned via the Confidence Network in each modality pathway. The confidence-weighted prediction is computed as:
@@ 3.4 Decision Making
The Mistake Correction Block serves as a refinement mechanism that further adjusts the final speaker prediction to correct errors. To this end, the four logit vectors are concatenated into a single input vector, and this high-dimensional vector is passed through a lightweight correction network,
@@
The weight of 0.2 serves as a scaling factor to prevent corrections from overly distorting the ensemble output.
Training follows a standard mini-batch procedure with curriculum-based sampling and optional mixup augmentation to improve generalization. We employed a cosine learning rate schedule with warmup and AdamW [adam] as the optimizer.
The loss function is designed as a multi-task training strategy, where the model optimizes multiple objectives simultaneously. It consists of the sum of multiple Focal Loss [lin2017focal] terms applied to each prediction head, ensuring balanced learning across different modalities. Additionally, label smoothing is incorporated to prevent the model from becoming overconfident.
Rather than summing these loss terms equally, the model applies variance-based weighting, inspired by [kendall2018multi], allowing it to dynamically learn the relative importance of each loss term. This is achieved by introducing learnable log-variance terms $\sigma^{2}$, leading to the adaptive loss formulation:
$\mathcal{L}=\sum_{i}\frac{1}{2\sigma_{i}^{2}}\mathcal{L}_{i}+\log\sigma_{i}$ (13)
4 Datasets
In this section, we present our procedure for establishing the first benchmark for person recognition on the CANDOR dataset. In addition, we use VoxCeleb1 as a standard dataset for multimodal person identification.
4.1 CANDOR Dataset
The CANDOR dataset [reece2023CANDOR], can serve as a benchmark for automatic speaker recognition. Despite its high potential for multimodal speaker recognition system development, this dataset has not been widely utilized due to the lack of a standardized partitioning scheme for training and evaluation and the absence of properly segmented data. To address this gap, we have prepared and publicly released the list of segments.
Fig. 2: Distribution of Speakers Across Sessions in the CANDOR Dataset
Table 1: Number of utterances and speakers in each dataset split for VoxCeleb1 and CANDOR.

### v2

@@ Abstract
Person identification systems often rely on audio, visual, or behavioral cues, but real-world conditions frequently result in missing or degraded modalities. To address this challenge, we propose a multimodal person identification framework that utilizes gesture as a situational enhancer to supplement traditional modalities like voice and face. Our model employs a unified hybrid fusion strategy, integrating both feature-level and score-level information to maximize representational richness and decision accuracy. Specifically, it leverages multi-task learning to process modalities independently, followed by cross-attention and gated fusion mechanisms. Finally, a confidence-weighted strategy dynamically adapts to missing data, ensuring that our single classification head achieves optimal performance even in unimodal and bimodal scenarios. We evaluate our method on CANDOR, a newly introduced interview-based multimodal dataset, which we benchmark in this work for the first time. Our results demonstrate that the proposed trimodal system achieves 99.51% Top-1 accuracy on person identification tasks. In addition, we evaluate our model on the VoxCeleb1 dataset as a benchmark and reach 99.92% accuracy in bimodal mode, outperforming conventional approaches. Moreover, we show that our system maintains high accuracy even when one or two modalities are unavailable, making it a robust solution for real-world person recognition applications. The code and data for this work are publicly available  [^1] The link has been removed for anonymization reasons..
I Introduction
@@
Typical person identification systems rely heavily on a single modality and risk severe performance degradation whenever that modality fails. This challenge calls for an approach that can effectively leverage the strengths of multiple modalities while downplaying those that are missing or degraded [9, 10]. However, multimodal person identification, which combines cues from several modalities like face, gait, gesture, and voice, can deliver robust and accurate recognition if the modalities are integrated to provide complementary support. Utilizing typical fusion methods [9, 12], which concatenate all features, pose still some challenges in real-world conditions, such as one or more modalities frequently becoming unreliable due to occlusions, poor lighting, background noise, or short utterances [11, 15].
@@
In this work, we utilize gesture information as a situational enhancer to provide supplementary information for person identification, specifically addressing the performance decrements that multimodal systems encounter during modality loss. The proposed system introduces an adaptive fusion strategy that integrates face, gesture, and voice modalities within a unified architecture (more in Section III). The core concepts of our approach to handle modality loss are: confidence weighting, cross-modality bridging and multi-stage fusion with mistake correction.
@@
In confidence weighting each modality is equipped with a confidence network that quantifies the reliability of each modality. Cross-modality bridging is implemented via cross-attention [33], enabling the learnable fusion of a primary modality with multiple auxiliary modalities to produce a more discriminative representation vector. Finally, mistake-correction module rectifies overconfident errors by re-examining logit outputs from the face, gesture, and voice pathways alongside the fused ensemble. The result is a flexible pipeline capable of robust person identification even in substantial modality loss.
@@
To validate our approach, we conducted experiments on the CANDOR dataset [30], providing a new benchmark for this recently introduced interview-style corpus. Furthermore, we used the audiovisual VoxCeleb1 dataset [24] as a benchmark to evaluate our approach compared to previous work. To this end, we utilized the original training, validation, and test splits provided by the dataset. It is essential to note that the multimodal version of VoxCeleb1 consists only of still images, which prevents capturing gesture modalities.
@@
In the following sections, we delve deeper into the specifics of our model and experimental findings. First, we discuss related prior work on modality fusion and highlight how existing methods work with different modalities. Next, we present the proposed trimodal system architecture. We then detail the partitioning of the CANDOR dataset and our data augmentation. Finally, our experimental results and ablation analyses illustrate the robustness of the proposed framework under unimodal, bimodal, and trimodal settings.
II Previous Work
@@
Initial face-gait fusion methods generally fused outputs at the score or decision level [13]. Some other early works relied on hand-crafted features and static fusion weights. For instance, Geng et al. proposed an adaptive scheme to combine face and gait cues in video, dynamically adjusting each modality’s contribution [13].
@@
Feature-level fusion then emerged, with approaches like subspace projection or concatenation to form a unified representation of face and gait [34]. While these multi-level fusion frameworks improved accuracy, they still relied on engineered features, limiting their performance in complex real-world scenarios.
@@
Studies commonly employ parallel pipelines for each modality, fusing their outputs at the feature or score level [28, 20]. Aung et al. employed transfer learning with Convolutional Neural Network-based (CNN) for both face and gait, concatenating feature vectors into a joint embedding and attaining above 97% accuracy [2]. Similarly, Manssor et al. tackled night-time surveillance by integrating face and gait recognition within a YOLO-based pipeline, highlighting the advantage of multimodal cues in adverse conditions [21].
@@
Recent research focuses on attention-based and context-aware fusion, leveraging transformer-like mechanisms to dynamically weight each modality. Prakash et al. introduced Adapt-FuseNet with a keyless attention module that assigns importance to face or gait features based on reliability, achieving improved identification performance in video-based tasks [26]. Zou and Wu proposed a robust hybrid framework with a dynamic weighting and distillation module that aligns and fuses face-gait features, improving recognition even under modality degradation [36]. These methods exemplify a shift from static to data-driven adaptive fusion, critical in unconstrained environments.
@@
While most deep fusion models rely on extensive labeled training data, emerging self-supervised techniques aim to reduce annotation dependency [4]. Such methods exploit natural co-occurrence of face and gait in videos, using tasks like contrastive learning or cross-modal reconstruction to learn shared identity embeddings without explicit identity labels.
@@
Tiong et al. [32] introduced a Transformer-based approach called Flexible Biometrics Recognition to fuse face and periocular features via dedicated attention modules and improved cross-modal interaction. Praveen and Alam [27] addressed face-voice verification using recursive joint cross-attention. By applying cross-attention repeatedly, the model aligns voice and face embeddings at multiple stages, progressively refining a shared audio-visual representation.
@@
Figure 1: Proposed Trimodal system. Consists of feeding each modality into its corresponding encoder and processing pathway. Subsequently, the outputs from these pathways are utilized within a Trimodal cross-attention block. Finally, all information is integrated into the decision-making block.
@@
Li et al. [15] proposed a learnable multimodal tokenizer for RGB, infrared, sketches, and textual information that encodes each modality into a shared embedding space, which is then fed to a frozen transformer backbone. Crucially, the proposed model synthesized missing modalities during training, fostering robustness under incomplete inputs.
@@
Several works have reported speaker identification results on the VoxCeleb1 dataset, leveraging both supervised and self-supervised methods. Nagrani et al. proposed a cross-modal voice-to-face matching network and achieved 81.0% Top-1 accuracy on a shared speaker pool [23]. Yadav and Rai introduced a VGG-style CNN trained with Softmax and center loss, reaching 89.5% Top-1 accuracy [35]. Chung et al. proposed environment-adversarial training with a thin ResNet-34 regularized via confusion loss, achieving 89.0% Top-1 accuracy [6].
@@
Shah et al. designed a dual-stream architecture fusing VGGFace and VGGVox embeddings, which reached 97.2% Top-1 accuracy in the bimodal setting [31]. Niizumi et al. introduced a self-supervised masked modeling duo approach that attained 81.2% Top-1 accuracy without any labeled fine-tuning [25], while Cheng et al. proposed a pipeline for integrating Speech Anonymization and Identity Classification (SAIC) pipeline for joint speaker anonymization and classification, improving performance to 96.1% [5]. Anidjar et al. employed Wav2Vec2 embeddings combined with a lightweight CNN and aggressive data augmentation, resulting in 90.5% Top-1 accuracy [1].
III Trimodal Person Identification
The proposed framework seeks to characterize each input video using a trimodal representation consisting of face, voice, and gesture. As illustrated in Figure 1, the system initializes three modality-specific pathways using feature vectors extracted from efficient pre-trained encoders. Beginning with the facial pathway, we utilize YOLOv8 [29] to detect and extract faces from the video frames. To ensure optimal input, we employ MagFace [22] for quality assessment, selecting the highest-quality facial frame per video. Finally, an IR101-AdaFace model [8], pre-trained on 12 million images from the WebFace12M dataset is used to extract a 512-dimensional embedding.
For the gesture modality, we utilize TimeSformer [3] to extract 768-dimensional embeddings. We apply the model to 16 equally spaced frames from each video, following the standard input configuration used by TimeSformer for action recognition tasks. Since video lengths range from 4 to 16 seconds, the sampling interval varies between 250 milliseconds and 1 second. To complete the trimodal representation, a ResNet293 [18], pre-trained on the VoxBlink2 [18] and VoxCeleb2 [7] datasets, is used to extract a 256-dimensional voice embedding from each utterance.
III-A Modality-Specific Pathways
@@
Each modality undergoes two main transformations, consisting of a fully connected block and a self-attention Block. Let $SA_{\text{face}}$ be the face pathway self-attention. Then, for the face modality, we have:
@@
where $\mathbf{x}^{(\text{face})}\!\in\mathbb{R}^{512}$ is the output of two dense layers, and $c^{(\text{face})}\!\in(0,1)$ is a scalar confidence value obtained via a small feed-forward head and a Sigmoid $\sigma(\cdot)$ function. This design allows the network to both refine face features and dynamically estimate their reliability, facilitating confidence-weighted fusion in subsequent modules. Moreover, $\mathbf{p}^{(\text{face})}\in\mathbb{R}^{K}$ (with $K$ the number of classes) is logit vector produced by the face-specific classifier head. Analogous expressions hold for gesture- and voice-specific pathways.
III-B Trimodal Cross-Attention
After $\mathbf{z}^{(\text{face})}$, $\mathbf{z}^{(\text{gest})}$, and $\mathbf{z}^{(\text{voice})}$ are extracted, they are fed into a cross-attention module in which all three modalities are jointly refined.
This module bridges the information across the face, gesture, and voice streams, ensuring that each modality interacts with the others rather than being processed in isolation. This interaction is realized through a two-stage process comprising cross-modality projection and cross-attention processing. In the projection stage, each modality receives a projection of the other two modalities into its respective feature space to facilitate alignment.
In the cross-attention stage, each modality-specific pathway acts as the Query ($Q$), while the other two modalities serve as the Key ($K$) and Value ($V$). This configuration allows the target modality to adaptively attend to and aggregate relevant features from the auxiliary streams. For instance, the face cross-attention sub-block utilizes the face embedding as the $Q$ to draw complementary information from the gesture and voice embeddings, which function as the $K$ and $V$.
This strategy enhances discriminative power by learning complex cross-modality relationships. Moreover, it improves the system’s robustness; if a specific modality is weak or missing, the model can still infer identity information from the remaining modalities via this adaptive sharing mechanism.
For instance, the face cross sub-block aggregates gesture and voice embeddings into the face branch:
@@ III-B Trimodal Cross-Attention
III-C Fusion Classification
To achieve the final classification, the framework integrates the three modalities through gated feature fusion and confidence-weighted fusion.
The fusion gate acts as a dynamic weighting mechanism, selectively emphasizing useful information and filtering out irrelevant noise. It consists of feature concatenation that brings together face, gesture, and voice features into a unified vector $\mathbf{z}_{x}^{\text{concat}}$, then a learnable gating mechanism determines how much each feature dimension contributes.
@@ III-C Fusion Classification
where $W_{1},W_{2}$ are learnable weight matrices, $\phi(\cdot)$ is a non-linear activation function, $\sigma(\cdot)$ is the Sigmoid function.
@@
$\mathbf{z}^{\text{fused}}=\mathbf{z}_{x}^{\text{concat}}\odot\mathbf{g}.$ (6)
@@
However, the confidence-weighted fusion mechanism ensures that the contribution of each modality is dynamically adjusted based on confidence scores. Each modality has an associated confidence score ($c^{(\text{face})},c^{(\text{gest})},c^{(\text{voice})}$) which are learned via the confidence network in each modality pathway. The confidence-weighted prediction is computed as:
@@ III-D Decision Making
The mistake correction block presented in Figure 1 serves as a refinement mechanism that further adjusts the final speaker prediction to correct errors. To this end, the four logit vectors are concatenated into a single input vector, and this high-dimensional vector is passed through a lightweight correction network,
@@
Empirically, the weight for the mistake-correction output was set to 0.2 to prevent it from disproportionately influencing the final ensemble prediction.
The loss function is designed as a multi-task training strategy, where the model optimizes multiple objectives simultaneously. It consists of the sum of multiple Focal Loss [17] terms applied to each prediction head, ensuring balanced learning across different modalities.
Rather than summing these loss terms equally, the model applies variance-based weighting, inspired by [14], allowing it to dynamically learn the relative importance of each loss term. This is achieved by introducing learnable log-variance terms $\sigma^{2}$, leading to the adaptive loss formulation:
$\mathcal{L}=\sum_{i}\frac{1}{2\sigma_{i}^{2}}\mathcal{L}_{i}+\log\sigma_{i}.$ (12)
IV Datasets
In this section, we present our procedure for establishing the first benchmark for person identification on the CANDOR dataset. In addition, we use VoxCeleb1 as a standard dataset for multimodal person identification.
IV-A CANDOR Dataset
The CANDOR dataset [30], can serve as a new benchmark automatic speaker recognition. Despite its high potential for multimodal speaker recognition, this dataset has not been widely utilized due to the lack of a standardized partitioning scheme for training and evaluation and the absence of properly segmented data. To address this gap, we have prepared and publicly released the list of segments.
Figure 2: Distribution of Speakers Across Sessions in the CANDOR Dataset
TABLE I: Number of utterances and speakers in each dataset split for VoxCeleb1 and CANDOR.